### v2

# Jakub Poćwiardowski, Mateusz Modrzejewski

Abstract
This work introduces the $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ composer system, capable of generating complete, multi-minute musical compositions with complex structures in any time signature, in the MIDI domain from input descriptions in natural language. The system utilizes an autoregressive transformer language model to map natural language prompts to composition parameters in JSON format. The defined structure includes time signature, scales, chord progressions, and valence-arousal values, from which accompaniment, melody, bass, motif, and percussion tracks are created. We propose a genetic algorithm for the generation of melodic elements. The algorithm incorporates mutations with musical significance and a fitness function based on normal distribution and predefined musical feature values. The values adaptively evolve, influenced by emotional parameters and distinct playing styles. The system for generating percussion in any time signature utilises probabilistic methods, including Markov chains. Through both human and objective evaluations, we demonstrate that our music generation approach outperforms baselines on specific, musically meaningful metrics, offering a viable alternative to purely neural network-based systems.
1 Introduction
Musicians around the world are constantly searching for inspiration to create groundbreaking and unique music. There has been a lot of work in recent years concerning generating music with the usage of AI techniques, both in symbol and audio domain. The latest state-of-the-art methods for generating music from text focus on the audio domain with works such as OpenAI’s Jukebox [1], Google’s MusicLM [2] or Meta’s MusicGen [3]. For symbolic composition, there are not many works capable of generating music from text, as there is not much text-symbolic music data and symbolic music is much more difficult to describe in natural language, especially for non-musicians. Some of the works trying to accomplish that are MuseCoCo [4] and BUTTER [5]. The most advanced symbolic models without prompt conditioning involve usage of Transformers in works such as [6, 7], which are also used to compose multitrack music [8, 9].
All the current state-of-the-art methods use neural networks to effectively replicate human music composition abilities. However, they depend on broad datasets, leading to reliance on dominant structures like 4/4 meter and common chord progressions. Furthermore, text-to-music models often fail to respond to specific musical terms, and audio-based music generation models produce outputs that are hard to edit further.
In this work, we present $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$, a tool for generating full multi-track songs in symbolic format and their subsequent editing with further prompts without relying on dominant musical structures. The system uses an autoregressive transformer model to map the input text to generated song structure and attributes. Further, it utilizes genetic algorithms along with probabilistic methods for generating MIDI (Musical Instrument Digital Interface) tracks for specific sections.
2 Related Work
2.1 Large language models for music generation
The recent revolution of Large Language Models (LLMs) invoked experiments concerning their abilities in tasks they were not initially created for. Examples of LLM usage for music composition include ChatMusician [10] which is a fine-tuned Llama2 model to treat Music ABC notation as a language, and ComposerX [11] which uses a multi-agent approach and chain-of-thought, where distinct instances of GPT-4 take care of interpreting user inputs, creating melodies, harmonising, and choosing instruments, among others.
2.2 Genetic algorithms for music generation
Genetic or evolutionary algorithms optimize melodies based on specified criteria, such as user feedback on quality. An early, well-known example of this approach is GenJam [12] meant for generating jazz solos. Due to the time-consuming nature of human feedback, latter approaches focus on developing objective criteria for fitness functions. [13] proposes 21 qualities related to melody, pitch, tonality, rhythm, and repetitiveness for this purpose. [14] uses music theory and charts, while [15] employs Gaussian distribution to model the criteria of the cost function.
2.3 Conditioning music generation
Non-neural music generation conditioning is usually based on predefined rules and emotions. The emotions are usually represented either as discrete emotions or as a point on a valence-arousal plane. [16] presents a rule-based mapping of valence and arousal values on musical qualities such as beats-per-minute (BPM), scales or pitch range. On the other hand, [17] proposes rules and fuzzy logic to map 12 emotions on tempo and other qualities. The predefined decision tree for chord selection is based on the circle of fifths, while the melody is created for the chords using genetic algorithms.
3 Structure of $M^{6}(GPT)^{3}$
The architecture was planned in order to have a separate components, which can be tested and developed independently, yet together create an innovative system. The system’s pipeline involves the following steps:
  1. 1.
    Predicting composition’s structure and parameters from text using LLM,
  2. 2.
    Generating melodic and drum tracks based on the structure and parameters provided by the LLM,
  3. 3.
    Integrating generated tracks into a MIDI file.
The structure of the system is presented in fig. 1. Although there is no direct feedback to the LLM, the generated structure and previous prompts provide context that allows for further modification.
Figure 1: The system architecture illustrating all key components. The MIDI Generator receives composition parameters, which it forwards to individual track generators. These generators produce tracks (each of them independently, based on the same input describing song structure) that are combined and encoded into a single MIDI file.
3.1 Predicting structure from text
To generate song structure and parameters, a LLM is used. We chose the GPT family because of its promising music theory knowledge and ability to use it via openai API, making it accessible from any computer, contrary to locally hosted LLMs requiring high computing resources.
To obtain the structure in an appropriate form, the model is provided with precise instructions as a system prompt:
"You are a music composing system. The user will ask you about the song they want to generate and your task is to respond with output of JSON file and JSON file only (…)"
The complete JSON template [^1] Full template and system prompt available at https://jpocwiar.github.io/M6-GPT3-Composer-Demo/. contains:
  • Song name,
  • Song sections with details on BPM, time signatures, scales, tracks (instrument and technique), chord progressions, and repeats,
  • Section placement in the song, including valence-arousal points [^2] Valence: $\langle-1,1\rangle$, Arousal: $\langle 0,1\rangle$, chosen for LLM comprehension.,
  • Composer’s note explaining artistic choices.
The model is instructed about the available set of scales, instruments, and chord types with precise indications to stick to this set. Another instruction is to maintain a coherent structure of the whole song so that the instrumentation and style of consecutive sections do not vary too much from each other unless prompted to do so.
The language model allows for iterative composition and editing based on user prompts. Users specify changes, and the model will adjust the structure accordingly, whether altering exact specified parameters or adapting the parameters to broader descriptions, such as sophistication, mood, or specific composer’s style.
4 Evolutionary Melodic Track Generator
The melodic sections in the presented multitrack systems are generated with genetic algorithms. Using this method, three types of tracks are generated:
  1. 1.
    Melodies - the most unrestricted tracks, playing mostly in mid-range, but capable of reaching high and low notes.
  2. 2.
    Bass - playing in low frequencies, usually more restricted and more repetitive than melodies.
  3. 3.
    Motif - usually playing in high frequencies, being the most repetitive track, adding texture to the sound
4.1 Encoding and Genetic Operations
To encode musical notes for optimization with a genetic algorithm, we used the same approach as [15, 14]. Notes are represented as MIDI values from 0 to 127, with rests marked as -1 and note extensions as -2. This encoding allows for different note durations while preserving a consistent bar length, which is essential for a multitrack system with multiple sections.
The operations used in our genetic algorithm are:
  • Random initialization,
  • One-point crossover,
  • Tournament selection,
  • Mutations with musical significance,
  • Fitness function based on the normal distribution.
4.2 Mutations with Musical Significance
We use mutations with musical significance to introduce musical features into melodies, which by default consist only of notes with uniform, minimum length. The choice of musical mutations was inspired mainly by [15, 12]. table 1 shows exemplary mutations used in our systems with examples of their impact on mutated melody.
Mutation Effect Example sequence
Original - 81 58 46 58 46 -2 -2 61
Interval Creates a random interval of -12 to 12 semitones between two notes. 81 89 46 58 46 -2 -2 61
Transpose Shifts a sequence segment by a random interval in $\langle-12,12\rangle$ range. 81 58 46 61 49 -2 -2 64
Extend Extends a random note, shortening (or removing) the preceding. 81 58 46 58 46 -2 61 -2
Rest Converts random note to a rest or a rest to a note. 81 58 46 58 -1 -2 -2 61
Long note Changes a random series of notes into extensions (-2). 81 58 -2 -2 46 -2 -2 61
-2 to note Converts a random extensions (-2) into a random note. 81 58 46 58 46 -2 41 61
Length norm. Divides long notes in two or extends short notes [^3] We treat a quarter note as the ”normal” length. Notes longer than this are randomly split, while shorter ones are extended.. 81 58 46 58 46 46 -2 61
Sort Sorts a random slice of a sequence in a random direction. 46 46 58 58 81 -2 -2 61
Repeat 1 Randomly chooses a sequence and pastes it to another measure. 81 58 46 58 81 58 -2 61
Repeat 2 Randomly chooses a sequence and pastes it immediately after. 81 58 81 58 46 -2 -2 61
Table 1: Mutation with musical significance (here, to illustrate the mutations in a table, we treat a measure as a 4-element sequence instead of factual 16).
4.3 Fitness Function
When designing the fitness function, we adapted the approach from [15], which employs a Gaussian distribution. In our implementation, the distribution’s mean represents the target metric value. The melody’s fitness score is determined by the weighted sum of each metric’s distribution value at the point corresponding to the actual measured metric value of the melody. The metrics used for genetic algorithm optimization include: the mean percentage of unique notes and intervals per measure; percentages of dissonant intervals, intervals larger than an octave, notes in the scale, notes in the current chord and rests in a sequence; tonal range; mean percentage of unique note lengths per measure; average note pitch and its deviation; length of strong beat notes; melodic contour; off-beat notes; normalized average interval size; logarithm of average note length and its deviation; counts of consecutive intervals within (1,3) and short notes (eighth or shorter); length of repeated fragments; and the number of measures beginning with a root note. All measured values are normalized to the (0, 1) range. The formula for fitness function calculation is presented in eq. 1.
$f(x)=\sum_{i=1}^{n}w_{i}\exp\left(-\frac{(r_{i}(x)-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)$ (1)
where:
  • $x$ is the melodic sequence,
  • $n$ is the number of features being measured in the sequence,
  • $w_{i}$ is the weight for $i$-th feature,
  • $\sigma_{i}$ is the standard deviation value for $i$-th feature,
  • $\mu_{i}$ is the desired value for $i$-th feature,
  • $r_{i}$ is the measured value of $i$-th feature in the sequence.
To ensure harmonic coherence in a multitrack composition, we add another metric to assess dissonance between distinct tracks. Our scoring method, based on [18], addresses these issues and also accounts for rests in one of the tracks. We evaluate intervals between distinct tracks modulo 12 at each index using scores from table 2. The total score is computed as the average of these individual scores. Finally, we normalize the average score for the entire song $a_{e}$ using the $tanh$ function as detailed in eq. 2.
Interval (semitones) Score
0, 3, 4, 8, 9 8
5, 7 15
1, 2, 10, 11 -20
6 -30
Rest in one of the two tracks 10
Rest in both tracks 0
Table 2: Scores assigned to various intervals between tracks.
$\tanh\left(\frac{a_{e}}{10}\right).$ (2)
The scoring system yields a range from -30 to 15 before normalization. To ensure that the highest scores do not disproportionately influence the results, we used 10 as the denominator in the normalization function. This approach rewards a variety of pleasing interval combinations while modestly favoring the most harmonic ones. Conversely, dissonances incur a swift penalty.
4.4 Melodic Tracks Modes
To support the diverse characteristics of melodic sections and assist LLM decision-making, we define three melodic sections and 11 generation modes, which influence genetic algorithm parameters.
For the main melody, the system operates in two modes: Melody (conservative tonal range and dynamics) and Solo (greater freedom, featuring fast and short notes).
Bass tracks include four modes: Short riff (one-measure riff repeated and shifted to fit chord roots), Long riff (two-measure riff with similar behavior), Bassline (lowest chord roots with randomized transitions and mutations for emotion), and Repetitive bassline (like bassline, but with consistent mutations per bar).
Motifs offer five modes: Long motif (one measure, repeated and shifted), Opening motif (half-measure with rests in the second half), Closing motif (rests in the first half), Repeated motif (half-measure repeated), and Short repeated motif (quarter-note length, repeated and trimmed to fit the sequence).
These modes, combined with valence-arousal values, define the genetic algorithm parameters. The rules draw inspiration from works [19, 20, 16] and our own experiences. table 3 presents the fitness function metrics, their desired values for specific tracks, and the influence of emotions.
Metric Melody Solo Bass Motif EI V A
Mean percentage of unique notes per measure Low High Med High None - -
Mean percentage of unique intervals per measure Med High High High Low $\uparrow$ $\uparrow$
Percentage of dissonant intervals Low Low Low Low Med $\downarrow$ -
Percentage of intervals bigger than octave Zero Zero Zero Zero None - -
Percentage of notes in scale High High High High Low - $\downarrow$
Percentage of notes in current chord Med Med High High Low $\downarrow$ $\downarrow$
Pitch range Low High Med Med Med - $\uparrow$
Percentage of rests in a sequence Low Low - - Med $\downarrow$ $\downarrow$
Mean percentage of unique note lengths per measure Low Med - - None - -
Average note pitch Med Med Low Med Med $\uparrow$ $\uparrow$
Pitch deviation Med Med Med - Med - $\uparrow$
Length of strong beat notes High High High - Med - $\downarrow$
Melodic contour Med Med Med Med High $\uparrow$ -
Off-beat notes Low Low Low - None - -
Average interval size (normalized to one octave) Med Med Med - Med $\uparrow$ -
Logarithm of average note length Med Med Med - Med - $\downarrow$
Deviation of logarithms of note lengths Med Low Low - Med $\downarrow$ $\uparrow$
Amount of consecutive intervals in range (1,3) High High - Med None - -
Amount of consecutive short notes (eight or shorter) Med High - - Med - $\uparrow$
Length of repeated fragments Med Med Med - Low - $\uparrow$
Measures beginning with a root note - - Med Med Low - $\downarrow$
Table 3: Desired metric values for melody, solo, bass, and motif, with emotion impact (EI) on their values. Arrows in V (valence) and A (arousal) show direction of change (proportional or inversely proportional).
5 Percussion Generator
To generate songs in odd time signatures, Deep Learning models like Transformers may not be ideal. Our analysis of the well-known Groove dataset [21] revealed that 98.96% of drum loops are in 4/4 time. Given the relative simplicity and commonality of drum patterns in many songs, we opted for a rule-based system incorporating probability and Markov Chains.
5.1 Data Representation
To encode drum states at a given time, we employed a binary representation of drum components proposed in [22]. Here, we use 12 components of percussion, namely: closed hi-hat, open hi-hat, bass drum, snare, 5 toms, crash, ride, and bell. To present an example 100000000000 and 001000000000 represent the closed hi-hat and bass drum respectively, while 101000000000 represents playing these components simultaneously. In default, one state corresponds to a sixteenth note.
5.2 Bass Drum and Snare
For the bass drum and snare, we analyzed fundamental patterns from a variety of sources, including drum courses, publicly available resources, and drum-specific datasets. Based on this analysis, we manually constructed a table of probabilities for these components in time signatures built on quarter-note and eighth-note divisions, with the number of beats per bar ranging from 2 to 9 (e.g., time signatures from 2/4 to 9/8).
For time signatures not covered by the predefined patterns, we apply a recursive decomposition approach. For instance, a 13/8 time signature is broken down into 7/8 and 6/8 patterns, while a 25/4 signature is split into 13/4 and 12/4, and further into 7/4, 6/4, 6/4, and 6/4. For time signatures with shorter notes, such as 11/16, we use patterns for longer notes (11/8 in that case).
As previously mentioned, each drum state is encoded with a sixteenth-note interval. Due to this representation, we pad the patterns with silent states, as detailed in eq. 3. Specifically, time signatures based on quarter notes include three silent states for each beat, while those based on eighth notes include one silent state per beat, and so forth.
$I=\frac{16}{{M_{R}}}-1$ (3)
where:
  • $I$ is a number of zero states added after every beat,
  • $M_{R}$ is the type of note (lower numeral of the time signature).
Hi-hats accent the tempo and are placed on quarter, eighth, or sixteenth notes, with patterns varying by the section’s emotion. Open hi-hats occasionally replace the last closed hi-hat in a measure to mark transitions, influenced by the arousal parameter. Ride cymbals and bells typically play on quarter notes, offset by an eighth, with variations based on emotional cues. Crash cymbals primarily occur at the start of measures, rarely appearing elsewhere.
Drum fills use all five General MIDI toms and the snare, generated based on fill probability and length. Fills are constructed using Markov chains, where each state is a 5-bit sequence representing drum activations, ordered from lowest to highest tom.
After adding all previously mentioned drum components, we obtain a full list of 12-bit sequences. To enhance diversity and realism, several modifications are applied to the patterns. The bass drum and snare drum occurrences are repeated in the following states with a certain probability. For states where the sum of active bits representing elements played with hands (bits 1, 2, and 5–12) exceeds 2, bits in this range are randomly zeroed until the sum is reduced to 2. Additionally, all bits from 5 to 12 are randomly zeroed with a specified probability for each state.
5.3 Percussion Modes
To enable the drum system to be effectively driven by the LLM, we define three playing modes: Only Beat (no drum fills), Drum Solo (only drum fills), and Standard (drum fills depend on emotional values, with higher probability in the final measure of a section).
Additionally, we offer three drum kits using General MIDI sounds: a standard kit, an ethnic kit, and an orchestral kit. The generation process is identical across kits, but binary indices correspond to different components based on the selected kit.
6 Chord generator
Chord progressions are the foundation of the composition, defined for each section based on the language model’s responses. The chord track is generated in one of three modes. In Continuous mode, the chord notes are sustained throughout, creating a seamless background. In Repeated mode, the notes are played together with lengths and rests determined by the arousal parameter, with up to 40% repetitions randomly omitted. In Arpeggio mode, the notes are played sequentially up and down the scale, with durations ranging from a sixteenth to a half note depending on arousal.
The chord names provided by LLM do not specify the octaves in which the individual notes are to be played, nor the possible repetition of these notes in other octaves. For $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Composer, we use valence-arousal values to apply chord inversions and modify voicing sizes. We partially base the solution on [16], where the authors propose that at the minimum value of arousal, a chord should consist of two notes, and at the maximum - of eight notes. In the context of our multi-track system, it was decided to limit the maximum number of notes in a chord to six. To ensure that the system works well for chords with varying numbers of notes (e.g., seventh or ninth chords), the following adjustments are made: If the arousal is below 0.3 or the chord has more than 4 notes, a fifth of the chord is removed. If the arousal exceeds 0.7 and the chord has fewer than 5 notes, a note an octave above the root is added. If the arousal exceeds 0.9 and the chord has fewer than 6 notes, a note forming an octave and the chord’s fifth interval is added.
Concerning the octaves in which the chord notes are played, [16] proposes concentrating the notes around the C4 note for the minimum value of valence and C6 for the maximum value. Taking into account that for our system, chords serve only as an accompaniment for a multitrack piece, with melody playing in a similar pitch register, it was decided to shift the range of central note concentration to span from A2 to A4. Based on the value of valence, the center of gravity of the notes is determined. Then, for each chord, the appropriate overturns are made using optimization to ensure that the notes do not overlap and that their center of gravity is as close as possible to the desired location. The conditioning of the chords via valence and arousal is presented in table 4.
Valence State Pitch Register Arousal State Voicing Size
Min Centered on A2 Min 2 Notes
Max Centered on A4 Max 6 Notes
Table 4: Influence of Emotional Values on Chord Parameters.
7 Experiemental evaluation
7.1 Experiemental Setup
In our experiments, we employ the gpt-4-1106-preview model with a temperature of 0. For all generations we use population size of 256 with 100 generations, tournament size of 4, 0.3 mutation rate and 0.9 crossover rate.
7.2 Subjective Listening Test
To assess the quality of the music generated by our model, we conducted a listening test with 23 participants, who were recruited from people who have basic musical knowledge and understanding of basic musical concepts.
The listening test was divided into two segments. In the first segment (table 6), participants evaluated 15 compositions from 3 different systems. Each composition was rated on a 1-to-5 scale across five criteria: 1) richness and diversity, 2) memorability, 3) entertainment value, 4) emotional conveyance, and 5) inspirational quality. The comparison systems were ComposerX [11] and MMT [9]. We selected 5 "best examples" that were more than a minute in length from the websites of both systems, randomly mixed them with our model’s 5 best compositions, and presented them to participants for evaluation.
Surveyed qualities (1-5 scale)
RD M EV CE I Avg
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ (ours) 3.40 ± 0.18 3.05 ± 0.20 3.34 ± 0.20 3.31 ± 0.18 3.36 ± 0.21 3.29 ± 0.09
MMT [9] 3.35 ± 0.19 2.98 ± 0.20 3.03 ± 0.22 3.34 ± 0.19 3.30 ± 0.22 3.20 ± 0.09
ComposerX [11] 3.07 ± 0.20 2.90 ± 0.21 3.04 ± 0.22 2.94 ± 0.23 2.87 ± 0.25 2.97 ± 0.10
Table 5: Comparison against the baseline models (RD: Richness and diversity, M: Memorability, EV: Entertainment value, CE: Emotional conveyance, I: Inspirational quality, Avg: Average of all criteria). Mean values and 95% confidence intervals are reported. According to the listeners, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ outperformed the other evaluated models in 4 out of 5 criteria, with its most significant advantage being the entertainment value of its compositions.
In the second test (LABEL:tab:subjective-results-midicaps), we compared the music generated by our system with pieces from the recently released MIDICaps dataset [23], which currently is the only openly available large-scale MIDI dataset with text captions. We randomly selected 5 full-length pieces from the dataset and generated 5 corresponding full-length songs using our system based on the same descriptions [^4] Generated pieces can be heard on https://jpocwiar.github.io/M6-GPT3-Composer-Demo/.. The participants were then surveyed to determine whether both the dataset compositions and the $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ generations accurately matched their descriptions regarding tempo, instruments, mood, structure, and overall impression.
Surveyed qualities (1-5 scale)
RD M EV CE I Avg
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ (ours) 3.40 ± 0.18 3.05 ± 0.20 3.34 ± 0.20 3.31 ± 0.18 3.36 ± 0.21 3.29 ± 0.09
MMT [9] 3.35 ± 0.19 2.98 ± 0.20 3.03 ± 0.22 3.34 ± 0.19 3.30 ± 0.22 3.20 ± 0.09
ComposerX [11] 3.07 ± 0.20 2.90 ± 0.21 3.04 ± 0.22 2.94 ± 0.23 2.87 ± 0.25 2.97 ± 0.10
Table 6: Comparison against the baseline models (RD: Richness and diversity, M: Memorability, EV: Entertainment value, CE: Emotional conveyance, I: Inspirational quality, Avg: Average of all criteria). Mean values and 95% confidence intervals are reported. According to the listeners, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ outperformed the other evaluated models in 4 out of 5 criteria, with its most significant advantage being the entertainment value of its compositions.
7.3 Objective Evaluation
To further support the listening test, we follow [9] and evaluate pitch class entropy, scale consistency, and groove consistency using the MusPy framework [24, 25]. For our system, we assess these metrics in three different scenarios. The “Prompt-driven” scenario involves generating full songs from text using an LLM. The “Focused” scenario standardizes the structure with a 4/4 time signature, 120 bpm tempo, C-F-Am-F chord progression in C Major and middle emotional values across all generations. In the “Randomized” scenario, all parameters are random except for the chords, which belong to the same scale. We compare our results with those presented in [9] and display them in table 7. We find that the outputs of $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ have comparable musical traits to a reference set of real music.
Objective evaluation results
Pitch Class Entropy Scale Consistency (%) Groove Consistency (%)
Ground truth (reference music) $2.974\pm 0.018$ $92.26\pm 1.25$ $93.05\pm 1.00$
MMM [8] $2.884\pm 0.023$ $93.13\pm 0.49$ $91.90\pm 0.64$
REMI+ [26] $2.897\pm 0.019$ $93.12\pm 0.51$ $\mathbf{92.90\pm 0.49}$
MMT [9] $2.802\pm 0.025$ $94.74\pm 0.42$ $92.09\pm 0.49$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Prompt-driven (ours) $\mathbf{2.907}\pm 0.067$ $\mathbf{92.640\pm 2.11}$ $98.919\pm 0.15$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Focused (ours) $2.849\pm 0.010$ $98.804\pm 0.14$ $98.558\pm 0.01$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Randomized (ours) $2.886\pm 0.021$ $95.113\pm 0.94$ $98.890\pm 0.11$
Table 7: Comparison with the baseline models with objective metrics following [9]. Mean values and 95% confidence intervals are reported, with the closest values to the ground truth highlighted in bold. Across three established metrics, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ with LLM-conditioned song structure achieves the best performance in two out of three cases.
8 Conclusions
In this work, we introduce $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$, a method for generating music from text without relying on extensive music datasets. Both objective metrics and subjective studies show that this approach can produce realistic and inspiring compositions. We believe that as fully neural-based systems continue to evolve, there is an increasing need to emphasize optimization-based techniques to create innovative and creative tools. The combination of controllability and structural flexibility of our approach with neural advances such as style morphing has the potential to push the boundaries of music generation in the future.